OpenStudy (anonymous):

Find the slope for the values in the table below. x -1 0 1 2 y 5 3 1 -1 A. 1 B. -1 C. 2 D. -2

OpenStudy (anonymous):

all you have to do is choose two points from the table and then use the slope formula: y2-y1/x2-x1 let me give you an example: lets say you chose these 2 points from the table: (1,4) and (2,2) where 1 is x1 4 is y1 2 is x2 and 2 is y2 now you just apply the formula: 2-4/2-1= -2/1=-2 the slope of that line is -2

OpenStudy (anonymous):

for each increase by 1 of x, y decreases by 2, so your slope will be -2

OpenStudy (anonymous):

ya it will be -2